The Mansfield Company manufactures and sells two lines of fishing rods. During the most recent accounting period, the Pro lin and the Novice line sold 15,000 and 2,000 units, respectively. The company's most recent financial statements are shown below Sales Less cost of goods sold S 900.000 S 240,000 Uni-level production cost Depreciation. production equipment 600,000 135000 50.000 125,000 Gross margin Less operating expenses: $ 175000 S 55000 Unit-level selling and admin Costs Corporate-evel facility expenses (fixed) 65.000 36,000 S 99.000 $(46.000) 40,000 36,000 Net income (loss) Based on this information, the company should a) Eliminate the Novice line because it is operating at a loss. b) Keep the forvice ine becaus it contributes $40,000 to total profitability c) Keep the Xovice line because it contributes $55,000 to toral profitability d) It is impossible to detemine with the given information.
osport.edu/dz/le/content/22310/viewContent/587167 Niew am 00:00 Juanita Merchant: Attempt 1 Time Taken:1:42:07 Question 20 (3 points) The Russell Company provides the following standard cost data per unit of product 3 Direct material (3 galons @ S6 per gallon) S 18.00 Direct labor (2 hours @S10 per hour) S 20.00 During the period, the company produced and sold 22,000 units incurring the following costs: Direct material 68000 galons @ $5.90 per galon Drect labor 4s.500hors @ $975 per hour 12 The direct labor usage variance was 15 18 a) $15,000 unfavorable b) s15,000 favorable C) $14,625 unfavorable d) $14,625 favorable 21 24
Answer #1

A) Eliminate the Novice line because it is operating at a loss

B) Labor usage variance = (Actual usage - Expected usage )*standard cost

So Labor usage variance = (45500-22000*2)*10 = 1500*10 = 15000

Since expected usage is more, so this is 15000 unfavorable